OpenStudy (anonymous):

y = -0.4x y = -0.8x + 2 I don't know how to do these AT ALL! Please help!!

OpenStudy (amistre64):

how to do what?

OpenStudy (anonymous):

well I don't know how to work this problem out

OpenStudy (anonymous):

(systems of equations)

OpenStudy (amistre64):

let y=y is a good start

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

we are given that: y = -0.4x and y = -0.8x + 2 so let y=y -0.4x = -0.8x + 2, solve for x

OpenStudy (anonymous):

so first do we add up 0.4 and -0.8?

OpenStudy (amistre64):

essentially yes, in a more mathical way we might want to say; lets add 0.4 to each side of the equals sign

OpenStudy (anonymous):

ok so now we have y=1.2+2?

OpenStudy (amistre64):

not quite .... -0.4x = -0.8x + 2 +0.4x +0.4x ---------------- 0 = -0.4x + 2

OpenStudy (anonymous):

oh ok so now we divide both sides by 2?

OpenStudy (amistre64):

while dividing by 2 is fine ... lets assume you want to subtract by 2 instead

OpenStudy (amistre64):

0 = -0.4x + 2 -2 - 2 -------------- -2 = -0.4x

OpenStudy (amistre64):

if we multiply thru by 10, it may help see the problem without the decimal -20 = -4x

OpenStudy (anonymous):

ok but, why are we multiplying by 10?

OpenStudy (amistre64):

because it may be helpful to see this wihtout the decimal point cluttering the issue. -20 = -4x is mentally a simpler thing to see than -2 = -0.4x

OpenStudy (anonymous):

oh I see so now we divide both sides by 4 and get 5 so x=5?

OpenStudy (amistre64):

correct :)

OpenStudy (anonymous):

ok what about solving for y?

OpenStudy (amistre64):

well, we know x=5 ... and they gave 2 equations that define y .... so use the definition of y that you like best, and set x to 5

OpenStudy (anonymous):

sorry I don't get that part

OpenStudy (amistre64):

y = -0.4x ... we know the value that x has to be sooo y = -0.4(5)

OpenStudy (anonymous):

ok Thank you!! :)

OpenStudy (amistre64):

youre welcome